OpenStudy (anonymous):

Auditions for a play in New York were held with 120 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many women auditioned?

OpenStudy (anonymous):

let x be the no. of men let y be the no. of women 120 = x + y 2x-3y = 20

OpenStudy (anonymous):

so work the rest on your own :) lol

OpenStudy (anonymous):

ughhh you and making me do the rest of the work! lol okay hold up ill try

OpenStudy (anonymous):

Okay! :D

OpenStudy (anonymous):

Ok here's a clue,, This is how you get how many men auditioned m=men w=Women m+w=120 3w-2m=20 2m+2w=240 3w-2m=20 5w=260 w=260/5 m+w=120 ---> m+52=120 ----> m=68 ... 68 me

OpenStudy (anonymous):

sooo w=52 ... 52 women

OpenStudy (anonymous):

thats not an answer on the thing! :o

OpenStudy (anonymous):

do you have answer choices

OpenStudy (anonymous):

44 40 76 80 yep there they are

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

lemme get some help

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

so what u wanna ask???

OpenStudy (anonymous):

whats the answer to this question

OpenStudy (anonymous):

I got what i though was the answer but, its not on his answer choices :(

OpenStudy (anonymous):

ok lemme try

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

oh alright so you dont know it?

OpenStudy (anonymous):

he's going to try!

OpenStudy (anonymous):

let x be the no. of men let y be the no. of women 120 = x + y 2x-3y = 20 this will help i think

OpenStudy (anonymous):

this is hard.. Lol

OpenStudy (anonymous):

yeah ill just skip it for now i guess... do you have a good guess on one of the answers?

OpenStudy (anonymous):

y=44

OpenStudy (anonymous):

i mean women are 44